A medication order calls for a dopamine drip at 5 mg/kg/min for a 185-lb patient. T he pharmacy adds 400 mg dopamine in 250 mL of D5W. Calculate the drip rates per minute when using administration sets delivering (a) 15 drops/mL, (b) 20 drops/mL, and (c) 60 drops/mL.
Answer #1

Dopamine calculation:  

mg/kg/min x wt in kg x 60

     mg/ml x 1000

mg/kg/min = 5

Weight in kg = 185/2.2 =84 kg

Mg/ml = 400/250 = 1.6

5x84x60 = 15.75ml/hr

1.6x1000

Drip rate with 15drops/ml set

ml/hr X drop factor    = 15.75 X 15/60 = 3-4 gtts/mt

total time in minutes

Drip rate with 20drops/ml set = 15.75 X 20/60 = 5 gtts/mt

Drip rate with 60drops/ml set = 15.75 X 60/60 = 15-16 gtts/mt